OpenStudy (anonymous):

Sandy invested $2,500 in a corporate bond that pays 9% interest compounded continuously. How long will it take for Sandy's investment to triple? Which equation correctly shows the information from the problem substituted into the equation?

OpenStudy (anonymous):

A. 2500=7500e^(.09t) B. 7500=2500e^(9t) C. 7500=2500e^(.09t) D. 2500=7500e^(9t)

OpenStudy (anonymous):

Is the answer C?

OpenStudy (anonymous):

\[A=Pe ^{decimal x Time}\]

OpenStudy (anonymous):

so yes! good job

OpenStudy (anonymous):

Thank you! So now it says "Divide both sides of the equation by 2,500. What is the result?" So far I have 3=e^(.09t)

OpenStudy (anonymous):

Wait so I got the right answer: 3=e^(.09t). Now it says to rewrite the exponential equation 3=e^(.09t) in logarithmic form. @RaeDeSol

OpenStudy (anonymous):

okay. so...... have you worked with natural logs before?

OpenStudy (anonymous):

I think so but I don't even know how to plug them in my calculator :/

OpenStudy (anonymous):

thats okay because we dont need a calculator to put something in logarithmic form.

OpenStudy (anonymous):

Oh alright then!

OpenStudy (anonymous):

let me write it out a take a picture because i cant work that equation button on here

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

when it says logarithmic form, does your teacher want the variable alone or what?

OpenStudy (anonymous):

I'm not sure... sorry... In the next question thought it says Use your calculator to solve the equation Ln3=.09t. soooo is that in logarithmic form?

OpenStudy (anonymous):

so then to solve for t..... (ln(3))/.09

OpenStudy (anonymous):

what does your calculator look like... maybe i can help you figure out how to use it

OpenStudy (anonymous):

Is the correct answer 12.21?

OpenStudy (anonymous):

YES! Great.

OpenStudy (anonymous):

Thank you so much!!!!

OpenStudy (anonymous):

Your welcome. you hardly needed me though